The sum of two numbers is 68. One number is 22 more than the other. Find the numbers.
Answer by TimothyLamb(4379)

x + y = 68
x = y + 22
---
put the system of linear equations into standard form
---
x + y = 68
x - y = 22
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 45
y = 23
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php